OpenStudy (anonymous):

Find x, y, and z. 9x-7y-9z= -73 18x-7y-9z= -91 -9x+7y+9z= -71

OpenStudy (wolf1728):

A) 9x-7y-9z= -73 B) 18x-7y-9z= -91 C) -9x+7y+9z= -71 Add equations A and C 9x-7y-9z= -73 -9x+7y+9z= -71 0x + 0y + 0z = -144 I'm thinking that since these 2 equations are identical (multiply equation A by minus1) there is no solution for this system of equations. (OR there are an infinite number of solutions).

OpenStudy (anonymous):

wolf...I think your right.